What is the price of gasoline per litre in Canadian dollars if a U.S gallon of gasoline costs us$3.74 one dollar is worth C$0.9739 and one U.S
gallon is equivalent to 3.8 litres

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

What is the price of gasoline per litre in Canadian dollars
A U.S gallon of gasoline costs us$3.74 one dollar is worth C$0.9739
us gallon = us $3.74 = = 3.74/0.9739 = C3.84
-------
and one U.S gallon is equivalent to 3.8 litres
Ans:: us gallon = C3.84/3.8 litres = C1.01 per litre
-------------------
Cheers,
Stan H.
-------------



Answer by Theo(13342)

the us dollar is currently worth 1.30 canadian dollars today..

your exchange rate is not valid today.

it probably was valid a few years ago, but the canadian dollar has lost value in comparison to the us dollar since.

however, we'll use it to solve your problem, even though it's not what the exchange rate is today.

we have:

1 us dollar is equal to .9739 canadian dollars.

1 gallon of gas costs 3.74 in us dollars.

that same gallon of gas would cost 3.74 * .9739 = 3.64 in canadian dollars.

1 gallon is equal to 3.8 liters.

therefore 3.8 liters = 3.64 in canadian dollars.

cost for 1 liter would therefore be 3.64 / 3.8 = .96 in canadian dollars.